OpenStudy (anonymous):

How do i calculate sin 120 without a calculator ?

OpenStudy (anonymous):

The answer is .580611184212314

OpenStudy (anonymous):

no like in ratio from

OpenStudy (anonymous):

it's 3 sqrt over 2 but idk how you get that :S

OpenStudy (anonymous):

I used a calculator

OpenStudy (anonymous):

Firstly find the square root of 3 from your calculator.. can you do this??

OpenStudy (anonymous):

sin(180-60) =sin180cos60-cos180sin60 =0-(-1)(3^.5/2) = sqrt(3)/2

OpenStudy (anonymous):

Then divide the result by 2 you will get your answer..

OpenStudy (anonymous):

in ratio format like how @nishika17 is doing it

OpenStudy (anonymous):

using the special triangles

OpenStudy (anonymous):

Oh do you want to know how sin120 = root(3)/2 ???

OpenStudy (anonymous):

yesss ! i dont really understand how

OpenStudy (anonymous):

sin(180-60) =sin180cos60-cos180sin60 =0-(-1)(3^.5/2) = sqrt(3)/2